I am having some trouble with this problem.
The length of a rectangle is twice the width. The area is 392yd^2. Find the length and the width.
The length is .........yd.
The width is .........yd.
Thank You

Answer by checkley71(8403)

L=2W
AREA=L*W
392=2W*W
2W^2=392
W^2=196
W=SWRT196
W=14 IS THE WIDTH.
L=2*14
L=28 IS THE LENGTH.
PROOF
14*28=392
392=393